OpenStudy (anonymous):

what is the slope intercet form of a line passing through (-6,1) and perpendicular to line y=-3x+1

OpenStudy (aroub):

y=mx+b; (-6,1), m= 1/3 1=1/3(-6)+b 1=-2+b 3=b => y=1/3x+3